OpenStudy (anonymous):

use the method of substitution to solve the sysem of linear equations.... 3x-y=3 7x-2y=9

OpenStudy (anonymous):

or using elimination method,

OpenStudy (anonymous):

no, its sustitution

terenzreignz (terenzreignz):

Well, substitution is easier to illustrate, anyhow :D From the first equation 3x - y = 3 Make it so that y is on one side of the equation, and everything else is on the other.

OpenStudy (anonymous):

let me see what you get,

OpenStudy (anonymous):

ill guide you there

OpenStudy (anonymous):

3x-y=3 (1) solve for y here and sub it to eq 2 7x-2y=9 (2)